OpenStudy (anonymous):

What is the slope of a line parallel to 4x + y = -1?

OpenStudy (raffle_snaffle):

Do you know how to take derivatives?

OpenStudy (anonymous):

u think if she is asking for th slope of y=mx+b SHE WOU KNOW DERIVATIVES ahhaha thts a good one

OpenStudy (raffle_snaffle):

I am thinking of this  it's parallel to the equation

OpenStudy (anonymous):

Yeah no I dont even know what a derative is actually.

OpenStudy (raffle_snaffle):

this is all you donirv

OpenStudy (anonymous):

Right i need the number though, and idk how to do it.

OpenStudy (anonymous):

make sence but if she knows derivatives she would hve problem with an equation in finding the slove of y=mx + b and i see ur point anyways lets help her IN THE EASY WAY * i dont mean derivatives way even thou is the faster way to do it*

OpenStudy (anonymous):

It gives me that equation, then wants numbers as the answer. Or rather one number.

OpenStudy (anonymous):

ok then u know the meaning of the word parallel right.. tht means tht 2 lines wou hve the sameeeeee.... wht??

OpenStudy (anonymous):

Yeah. I know that.

OpenStudy (anonymous):

same slope right..!! so lets find the slope of this equation then 4x + y = -1? lets leave Y alone to change it in the standard form y= -4x -1 y= mx+b m is slove in this case -4 wou be the slove of our parallel line

OpenStudy (anonymous):

do u understand?

OpenStudy (anonymous):

So I find the slope of the equation, and its that number?

OpenStudy (anonymous):

yes cuz the word parallel means tht they hve the same slove like this..